OpenStudy (anonymous):

A man is currently 20 yrs older than his oldest son. His oldest son is currently 2 yrs older than his youngest son. 3 yrs from the current time, the product of ages of his sons will be ten more than 11 times the age of the father. What is the current age of mans youngest son?

OpenStudy (anonymous):

you sould create a variable for the youngest son's age and write equations for the given info in terms of that variable and then solve the resulting system for that variable

OpenStudy (anonymous):

ok so let Youngest son age = x Oldest = x+2 Dad = x + 2 + 20 = x + 22 3 years from now: Youngest = x+3 Oldest = x + 2 +3 = x+5 Dad = x + 22 + 3 = x + 25 (x+3)(x+5) = 11(x+25) + 10 x^2 + 8x + 15 = 11x + 275 + 10 X^2 + 8x - 11x + 15 - 285 = 0 x^2 -3x -270 = 0 Factor: (x-18)(x+15) = 0 x = 18 or -15....age can't be negative so x = 18 Current age was x for youngest. Youngest = x = 18 years Also: Oldest = x+2 = 20 Dad = x+22 = 18 + 22 = 40 years @burrell

OpenStudy (anonymous):

hope it helps!! )